step by step how would u do a problem like 1/2y-2/3=7/8
Answer by checkley71(8403)

1/2y-2/3=7/8
1/2y=7/8+2/3
1/2y=(7*3+2*8)/24
1/2y=(21+16)/24
1/2y=37/24
24y=37*2
24y=74
y=74/24
y=37/12
y=3 1/12